OpenStudy (anonymous):

If the revenue of a shoe company is f(x) = 33x + 16, where x is the number of pairs of shoes sold, and the expenses of the company are g(x) = 4x + 50, where x is also the number of pairs of shoes sold, what is the profit of the company if they sell 105 pairs of shoes? A. $3,481 B. $3,011 C. $3,951 D. $3,111

OpenStudy (mathmate):

Profit = f(x)-g(x) = f(105)-g(105)

OpenStudy (anonymous):

what that means i stiill dont get it

OpenStudy (mathmate):

f(105) is a function notation which means that from the given definition of f(x)=33x + 16, then f(105) would evaluate the right-hand side replacing all occurrences of x by 105, giving f(105)=33(105)+16=3465+16=3481 You can evaluate g(x) in a similar way. Since profit=revenue - expenses, so profit(x)=f(x)-g(x) where x is the number of items sold.